OpenStudy (anonymous):

Miguel is designing shipping boxes that are rectangular prisms. The shape of one box, with height h in feet, has a volume defined by the function V(h) = h(h – 5)(h – 6). Graph the function. What is the maximum volume for the domain 0 < h < 6? Round to the nearest cubic foot.

OpenStudy (anonymous):

How would I solve this?

OpenStudy (anonymous):

Have you tried graphing it? Or attempting it?

OpenStudy (anonymous):

I'm not sure HOW to graph it. e_e

OpenStudy (anonymous):

V(h) = h(h – 5)(h – 6) is of the form of a factored cubic To graph it, set h(h – 5)(h – 6) = 0 What is h equal to just by looking at the equation?

OpenStudy (anonymous):

@xKingx

OpenStudy (anonymous):

Srry slow pc, but to answer your earlier question, I'm not sure

OpenStudy (anonymous):

You know that if a x b x c = 0 Then ONE of the variables (a or b or c) MUST be equal to 0. Do you understand this

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Then if h(h – 5)(h – 6) = 0 h = 0 h-5 = 0 h-6 = 0

OpenStudy (anonymous):

Do you understand this

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

What is h equal to then

OpenStudy (anonymous):

0

OpenStudy (anonymous):

So it would go 0<0<6, but that wouldn't give me any cubic feet to round.

OpenStudy (anonymous):

Nope You forgot h-5=0 and h-6 =0 From h-5=0, we can see that h=5

OpenStudy (anonymous):

So what is h equal to?